### v2

# Taming Latent Diffusion Models to See in the Dark

Abstract
Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image is a significant challenge in computational photography. Due to the limitation of large-scale paired data, prior approaches have difficulty in recovering fine details and true colors in extremely low-light regions. Meanwhile, recent advancements in generative diffusion models have shown promising generating capabilities, which inspires this work to explore generative priors from a diffusion model trained on a large-scale open-domain dataset to benefit the low-light image enhancement (LLIE) task. Based on this intention, we propose a novel diffusion-model-based LLIE method, dubbed LDM-SID. LDM-SID aims at inserting a set of proposed taming modules into a frozen pre-trained diffusion model to steer its generating process. Specifically, the taming module fed with low-light information serves to output a pair of affine transformation parameters to modulate the intermediate feature in the diffusion model. Additionally, based on the observation of dedicated generative priors across different portions of the diffusion model, we propose to apply 2D discrete wavelet transforms on the input RAW image, resulting in dividing the LLIE task into two essential parts: low-frequency content generation and high-frequency detail maintenance. This enables us to skillfully tame the diffusion model for optimized structural generation and detail enhancement. Extensive experiments demonstrate the proposed method not only achieves state-of-the-art performance in quantitative evaluations but also shows significant superiority in visual comparisons. These findings highlight the effectiveness of leveraging a pre-trained diffusion model as a generative prior to the LLIE task. Project page: https://csqiangwen.github.io/projects/ldm-sid/.
{strip}
Figure 1: An extremely low-light real-world image from SID Dataset [4] with a ×$\times 300$ amplification ratio. The image processed by our proposed method exhibits clear improvement in structural content, showcasing finer details compared to previous works. The description generated for each image by InstructBLIP [6] also suggests that our result provides higher quality to downstream tasks.
1 Introduction
Low-light image enhancement (LLIE) is a crucial task in computational photography, particularly for RAW images. Enhancing a low-light noisy RAW image into a well-exposed and clean sRGB image not only requires addressing noise reduction but also grapples with challenges related to color correction, among other image signal processing operations.
In recent years, learning-based methods [4, 14, 38, 32, 30, 37, 5, 12, 43, 11] have emerged as the dominant approach and have made remarkable progress in the LLIE task. Several studies [4, 38, 22, 5, 12, 34, 31, 36] collect datasets with real-world noisy-clean data pairs for training these deep models. Nevertheless, due to the limited range of scenes that these datasets cover, learning-based methods inevitably suffer from noticeable artifacts, e.g., over-smooth and color distortion problems, in extremely dark regions.
An intuitional way to improve the performance of learning-based methods is to increase the number of paired training data. To avoid the challenges associated with laborious real-world data collection, some works [9, 32, 28, 1, 3, 38] focus on designing noise models to synthesize realistic noisy-clean data pairs to train LLIE methods on much more data which improves the generalization. However, the gap between the synthetic and real-world noise distribution still impedes LLIE methods to achieving satisfying restoration and denoising performance in challenging cases as mentioned in [41, 17, 16].
With the rapid development of generative models, especially diffusion models, some works [8, 27, 39] demonstrate the generative prior from a pre-trained diffusion model shows promising performance in image restoration tasks. Instead of burdensomely training a diffusion model from scratch [42, 30, 18, 13] on a specific dataset, a diffusion model trained on a large-scale open-domain dataset still shows powerful generating ability in a target task.
Inspired by recent works [8, 27, 39], we propose a novel LLIE method, LDM-SID, based on a frozen pre-trained diffusion model [20]. To fully exploit the generative prior, instead of directly finetuning the parameters of the diffusion model, LDM-SID aims at inserting a set of trainable taming modules, whose key part is the spatial feature transformation (SFT) layer [29] to modulate intermediate features of the diffusion model for the LLIE task. Meanwhile, with an observation that there are dedicated generative priors in different portions of the diffusion model, we propose to employ 2D discrete wavelet transforms to divide the LLIE task into two parts: low-frequency content generation and high-frequency detail maintenance. To be specific, LDM-SID employs 2D discrete wavelet transforms on the low-light noisy RAW image to extract its low-frequency and high-frequency sub-bands. Then, the low-frequency sub-band is fed into taming modules to steer the structural content generation while the high-frequency is adopted to maintain the precise and clear details in the final sRGB output. Feeding low- and high-frequency subbands separately allows us to skillfully exploit dedicated generate priors in different portions of the diffusion model. By training the proposed taming modules on a small-scale LLIE dataset, LDM-SID finds an optimized balance between image fidelity and perceptual quality, which allows it to achieve state-of-the-art LLIE performance by generating high-quality and artifact-free images.
To sum up, our primary contributions are:
  • We propose a novel diffusion-model-based LLIE method, LDM-SID, which is simple but effective. Instead of training a diffusion model on a specific LLIE dataset, LDM-SID only needs to insert a set of trainable taming modules into a pre-trained diffusion model to generate high-quality outputs. Keeping the pre-trained diffusion model frozen allows LDM-SID to fully exploit generative priors from the diffusion model trained on a large-scale open-domain dataset.
  • To maximally exploit dedicated generative priors in different portions of the diffusion model, we further propose to divide the LLIE task into low-frequency content generation and high-frequency detail maintenance by applying 2D discrete wavelet transforms (DWT) on the input RAW image. This ensures that LDM-SID achieves an optimized balance between generating accurate structural content and achieving fine-grained detail enhancement with noise reduction, resulting in improved image quality.
  • Extensive evaluations conducted on four widely-used real-world datasets show the proposed method not only achieves state-of-the-art performance according to quantitative metrics but also significantly outperforms previous works in perceptual quality. Our LDM-SID excels in challenging regions by producing outputs without any remaining noise, highlighting its superiority over prior works.
Figure 2: The overview of our proposed LLIE method, LDM-SID. The Bayer RAW image is processed by similar operations mentioned in [4]. A series of 2D discrete wavelet transforms (DWT) is applied to the processed image for capturing the low-frequency (LL) and high-frequency subbands (LH, HL, HH). The low-frequency subband (LL) serves to modulate the feature at each layer in the LDM. Specifically, each feature has its corresponding taming module, whose key part is an SFT layer, to map the sub-band into a pair of scale $\gamma$ and shift $\beta$ parameters. Similar to the low-frequency taming, the features of the decoder $\mathcal{D}$ are modulated by another set of taming modules, where the LL sub-band is mapped to the scale $\gamma$ and the concatenation of LH, HL, HH is mapped to the shift $\beta$. All parameters from the pre-trained Stable Diffusion are frozen and only taming modules are trainable.
2 Related Work
2.1 Low-light Image Enhancement
For enhancing low-light RAW images, numerous studies have showcased the promising performance of deep learning-based methods. Several studies [4, 30, 37, 5, 12, 43, 11] propose supervised learning-based methods that leverage real-world paired data for training. Chen et al[4] propose the first RAW dataset, SID, which contains 5096 RAW data pairs and they designed a U-Net-based method to tackle the LLIE problem. Zhu et al[43] propose an edge-enhance multi-exposure fusion network to improve the performance of a learning-based method in the LLIE task. Wang et al[30] propose the ExposureDiffusion which offers an effective approach for reducing noise in low-light images by leveraging the principles of the diffusion model [10]. Although the supervised learning-based methods have shown promising performance, they still suffer from restoring background with noticeable artifacts in heavily noisy areas due to the limited range of scenes that the dataset covers. Since extending the scale of real-world datasets is quite challenging, diverse noise models [32, 38, 9, 28, 1, 3] are proposed to generate realistic noise for synthesizing more low/normal-light data pairs. Physics-based noise modeling [9, 32, 41] has been commonly used in previous works for noise synthesis in low-light RAW images. To improve the accuracy of the unknown noise synthesis, some works [1, 3, 38] focus on designing learning-based noise modeling. However, the gap between the synthetic and real-world noise distributions impedes learning-based methods to achieve satisfactory performance in real-world photography.
In this work, instead of training a method on a large-scale synthetic dataset to improve the restoration performance in the LLIE task, we propose to exploit generative priors from a powerful text-to-image diffusion model [20], which is pre-trained on a large-scale sRGB dataset, LAION-5B [21] containing 5.85 billion text-image pairs. With the help of our proposed taming modules, the proposed method LDM-SID solely trains a set of taming modules on the existing small-scale LLIE dataset (e.g., SID-Sony [4]) to tame the pre-trained diffusion model instead of finetuning any parameters in it.
2.2 Prior for Low-light Image Enhancement
Instead of relying on more data pairs to enhance the restoration performance, several studies [33, 8, 35] have focused on incorporating priors into their methods to improve the LLIE performance. Wu et al[33] propose to incorporate the semantic priors into the LLIE method. This method achieves more natural results compared to previous works by integrating image features with semantic features extracted from a pre-trained segmentation network. Xu et al[35] propose to take an SNR (Signal-to-Noise-Ratio) prior to their method to achieve better LLIE performance. Their method first calculates an SNR map from the input noisy image and its corresponding clean image which is generated by a no-learning-based denoising operation. The SNR map is fed into an SNR-aware transformer [25] to guide feature fusion. Most recently, Fei et al[8] propose to utilize the generative prior from a pre-trained diffusion model [7] for unified image restoration, including the LLIE task. By optimizing a carefully designed degradation model, their method regards the input low-light image as the guidance to control the generating process. Since the generated image is supervised by the input low-light image, the heavy noise from the input still remains in the final output. Although the above methods outperform previous works in the LLIE task, they still suffer from noticeable artifacts or remaining noise. Additionally, sRGB input is required by these methods to exploit corresponding priors, which impedes them from achieving LLIE in RAW space.
In this work, the proposed LDM-SID fully exploits the generative prior from a frozen pre-trained diffusion model [20] without any degradation assumption. The generative prior of the diffusion model trained on a large-scale dataset helps LDM-SID achieve noise-free LLIE performance.
3 Generative Priors in the Diffusion Model
Figure 3: A text-to-image example from Stable Diffusion. The difference map between the latent representation and sRGB output indicates that the latent representation generated by the LDM shows essential structural content while the decoder $\mathcal{D}$ introduces abundant details but scarcely modifies the global structure for the final sRGB output. The latent representation is converted to sRGB visualization by the linear transformation [15].
We utilize the Stable Diffusion [20] as our backbone. Stable Diffusion is a text-to-image generative model trained on the LAION-5B dataset [21] which is approximately 3M times bigger than a widely-used LLIE dataset SID-Sony [4]. It consists of an autoencoder and a latent diffusion model (LDM). Delving into the autoencoder, the encoder $\mathcal{E}$ serves to encode an sRGB image into a latent representation $Z$ while the decoder $\mathcal{D}$ reconstructs the sRGB image from $Z$. The LDM performs the denoising process for the noised latent representation $Z_{t}$, where $t$ denotes the timestep.
Recently, Wang et al[27] propose to exploit generative priors of Stable Diffuiosn to benefit the image super-resolution task. However, they do not delve into the presence of dedicated generative priors within different portions of Stable Diffusion. With the further study of these priors within Stable Diffusion, we have two observations shown in Fig. 3. 1) Given a text prompt, the LDM primarily serves to generate the latent representation which shows high-correlation structural information to the prompt and neglects to include intricate details; 2) During decoding the latent representation to the sRGB output, the decoder $\mathcal{D}$ introduces abundant details to improve the perceptual quality of the output but scarcely modifies the global structure. These facts indicate that the LDM mainly contains the low-frequency generative prior while the decoder $\mathcal{D}$ chiefly has the high-frequency generative prior.
4 Method
An overview of the proposed method is shown in Fig. 2. Given a low-light noisy RAW image, we first process it to match the illumination to our target sRGB image: $X=(\text{RAW}-\text{Black Level})\times\gamma$, where $\gamma$ represents the user-defined amplification ratio. We aim at taming Stable Diffusion to perform the image signal processing pipeline (ISP), restoration for over-dark areas, and denoising for the processed image.
Based on the above observation that the LDM is adept at generating structural content, which is predominantly dictated by low-frequency information, whereas the decoder demonstrates proficiency in the generation of high-frequency details as shown in Fig. 3, we are inspired to extract different-frequency sub-bands from the input image by 2D discrete wavelet transforms (DWT) and separately feed them into the LDM and decoder $\mathcal{D}$ to exploit dedicated generative priors.
4.1 Low-frequency Content Generation
Figure 4: The low-frequency sub-band, extracted using 2D discrete wavelet transforms (DWT), reveals clearer structural information compared to the low-light input. On the other hand, merely rescaling the input to align with the latent size of the LDM indicates that the information remains predominantly influenced by the noise. The result “Ours (Rescaling)” referring to merely rescaling the input for taming the LDM shows over-smooth and poorly-structured results due to the negative impact of the noise.
Employing a series of DWT not only extracts the low-frequency sub-band (LL) but also allows us to effectively align the input information with the latent size of the LDM. Compared to directly rescaling the input RAW image, feeding the low-frequency sub-band (LL) to the LDM enables the model to concentrate on generating high-correlation structural content while reducing the negative impact of the noise, as shown in Fig. 4.
By initializing a latent representation $Z_{T}\sim\mathcal{N}(0,I)$, we aim at steering the denoising process in the LDM by inserting a taming module at each layer. After denoising $T$ timesteps, we obtain a clean latent representation $Z_{0}$ showing a high correlation to the low-frequency sub-band (LL).
Taming Module
For the taming modules of the LDM, we employ a similar time-aware guidance strategy as discussed in [27]. However, instead of deriving the structural information by encoding the input image with the encoder $\mathcal{E}$ as in  [27], we directly utilize the low-frequency sub-band (LL) to provide our structural information. This liberates us from the substantial computational cost of encoding.
We insert a taming module at each layer in the LDM. The taming module consists of a set of convolutional layers and an SFT layer for mapping the low-frequency sub-band (LL) to a pair of affine transformation parameters:
[…]
where $l$ denotes the layer number in the LDM and t denotes the timestep. Then, the feature $\textbf{\text{F}}^{\text{LDM}}_{l}$ at the layer $l$ in the LDM is modulated by
$\hat{\textbf{\text{F}}}^{\text{LDM}}_{l}=(1+\alpha_{l})\odot\textbf{\text{F}}^{\text{LDM}}_{l}+\beta_{l}.$ (2)
After taming the LDM to generate the latent representation $Z_{0}$, an sRGB can be directly obtained by feeding it to the decoder $\mathcal{D}$. As shown in Fig. 7, such sRGB image presents well-exposed and high-correlation structural content but inaccurate color and over-generated details. This inspires us to tame the decoder $\mathcal{D}$ during upsampling to guarantee exact illumination and precise details in the final sRGB image.
4.2 High-frequency Detail Maintenance
This high-frequency “taming” aims to leverage the noise-free and fine-grained generative prior in the decoder to produce final sRGB images with enhanced details and devoided noise.
Similar to the taming process in the LDM, we insert a taming module at each upsampling layer in the decoder $\mathcal{D}$. For taming the decoder $\mathcal{D}$, not only the high-frequency subbands are needed but also the low-frequency. There are two key reasons for that:
  • To ensure the final sRGB image exhibits the correct illumination, the low-frequency sub-band (LL) is mapped to the scale parameter in the SFT layer. This strategy enables a linear adjustment of the intensity of decoder feature values during the modulation.
  • For steering the generation with precise details during upsampling, the high-frequency sub-bands ([LH, HL, HH]) are mapped to the shift parameter in the SFT layer for adding detailed information onto the adjusted decoder feature.
4.3 Training Objective
For training the taming modules of the LDM, the ground-truth sRGB image $Y$ is first encoded into a latent representation $Z_{0}$ via the encoder $\mathcal{E}$. Then, the latent representation is corrupted with noise using the standard forward diffusion schedule [10]: $Z_{0}\overset{\epsilon}{\rightarrow}Z_{t}$, where the noise $\epsilon\sim\mathcal{N}(0,I)$ and the timestep $t\in\left[1,T\right]$. Since the LDM serves to predict the noise from the noised representation, the training objective is
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{Z_{0},\epsilon\sim\mathcal{N}(0,I),t}\left[\left\|\epsilon-\epsilon_{\theta}(Z_{t},t,\tau^{\text{LDM}}_{\phi}(\text{LL}))\right\|^{2}_{2}\right],$ (3)
where $\epsilon_{\theta}(\cdot)$ denotes the LDM and $\tau^{\text{LDM}}_{\phi}(\cdot)$ denotes taming modules.
For training the taming modules of the decoder $\mathcal{D}$, a clean representation $\hat{Z_{0}}$ is generated by our tamed LDM for each low-light input image and the training objective is
$\mathcal{L}_{\mathcal{D}}=\mathbb{E}_{\hat{Z}_{0}}\left[\left|\mathcal{D}(\hat{Z}_{0},\tau^{\mathcal{D}}_{\phi}(\text{LL},\text{H}^{\ast}))-Y\right|\right],$ (4)
where $\text{H}^{\ast}$ denotes the concatenation of high-frequency subbands (LH, HL, HH) and $\tau^{\mathcal{D}}_{\phi}(\cdot)$ denotes taming modules.
Note that all parameters from the LDM and decoder $\mathcal{D}$ are frozen and only our proposed taming modules are trainable.
Figure 5: The qualitative evaluations on SID-Sony [4] (1-4 rows) and ELD-Sony [32] (5-6 rows) datasets (Zoom-in for best view). The input images are amplified and post-processed with the same operations as ELD [32]. The proposed method shows notable superiority in recovering structural information and enhancing details in extremely dark and noisy regions. In the $\text{5}^{\text{th}}$ row, the proposed method recovers the text much more effectively than previous works.
Table 1: The quantitative comparison in extremely challenging scenarios from different datasets. The best score is denoted as boldfaced. The proposed method significantly outperforms previous works according to perceptual metrics, including LPIPS [40], NIMA [24] and CLIP-IQA [26]. It indicates the sRGB results obtained from the proposed method exhibit significantly clearer global structure and more natural details compared to those from previous works.
Dataset
Amplification Ratio
Methods
PSNR $\uparrow$
SSIM $\uparrow$
LPIPS $\downarrow$
NIMA $\uparrow$
CLIP-IQA $\uparrow$
SID-Sony [4] $\times$300 SID [4] 27.1862 0.8219 0.3906 3.0481 0.2279
ELD [32] 27.4026 0.8271 0.3876 3.0346 0.2342
LED [14] 26.3541 0.8228 0.4168 3.1831 0.2688
ExposureDiff [30] 28.0544 0.8383 0.3892 2.9931 0.2519
Ours 27.2857 0.8113 0.2749 3.5583 0.3248
ELD-Sony [32] $\times$200 SID [4] 28.9735 0.7990 0.2618 2.9780 0.2057
ELD [32] 29.0467 0.8063 0.2585 3.1798 0.2504
LED [14] 30.4692 0.8769 0.2058 3.1718 0.2883
ExposureDiff [30] 30.7530 0.8582 0.1975 3.0221 0.2915
Ours 29.8284 0.8906 0.2035 3.4397 0.3448
LRD [38] -3EV LRD [38] 29.4312 0.8784 0.2198 3.1985 0.3078
Ours 29.0672 0.8889 0.1858 3.6180 0.3862
DRV [5] [$\times$100, $\times$300] Retinexformer [2] 29.3872 0.9030 0.1460 4.7620 0.3326
Ours 28.4360 0.8900 0.1304 4.7839 0.4318
Figure 6: The qualitative evaluations on LRD [38] (the first row) and DRV [5] (the second row) datasets (Zoom-in for best view).
5 Experiments
5.1 Implementation Details
The proposed method is evaluated on four widely-used datasets for the LLIE task, including SID-Sony [4], ELD-Sony [32], LRD [38], and DRV [5]. The train and test subsets follow the same strategy as previous works. Since the proposed method is fed with a RAW input and outputs an sRGB image, we convert the RAW output of previous works with the same ISP pipeline (e.g., rawpy [^1] https://pypi.org/project/rawpy/) to guarantee fair comparisons.
During training, the taming modules of the LDM are trained for 100 epochs while the taming modules of the decoder are trained for 30 epochs. By initially taming the LDM, it gains the capability to generate the clean latent representation $Z_{0}$ for each low-light training image. This generated latent representation is subsequently utilized to train the decoder taming modules. The training process is performed on 512 $\times$ 512 resolution with 4 NVIDIA GeForce RTX 3090 GPUs. During inference, the LDM follows the DDIM sampling strategy [23] with 200 timesteps.
5.2 Quantitative Evaluation
In this work, we mainly focus on the perceptual quality of results which are enhanced from extremely dark and noisy images. Besides distortion metrics (PSNR and SSIM), we also adopt LPIPS [40], NIMA [24], and CLIP-IQA [26] for our quantitative evaluations. LPIPS [40] and NIMA [24] are two widely-used reference-based and non-reference-based neural perceptual metrics. CLIP-IQA [26] is a state-of-the-art non-reference-based image quality perception assessment method. It is based on a large model, CLIP [19], to assess an image considering quality and abstract perceptions.
According to Table  1, the proposed method outperforms other methods in terms of non-reference-based perceptual metrics, including NIMA [24] and CLIP-IQA [26]. It indicates the sRGB results obtained from the proposed method exhibit significantly more natural details compared to those from previous works. The proposed method does not achieve the best but still competitive performance in terms of PSNR metric. There are two potential reasons: 1) Some misalignment exists between the low-light input and long-exposure ground truth; 2) Noticeable noise still remains in the long-exposure ground truth [38]. The structural quality assessment based on SSIM and LPIPS [40] indicates the proposed method has substantial advantages in enhancing low-frequency information, especially structural content.
Figure 7: The visualization of the ablation study. Different taming strategies (rescaling and DWT) and taming targets (LDM and decoder $\mathcal{D}$) are combined in this study. Taming the LDM by rescaling the input information tends to result in an ambiguous global structure. In contrast, the taming result using 2D discrete wavelet transforms (DWT) exhibits distinct structural content. Moreover, taming the decoder $\mathcal{D}$ results in improved color distribution and the enhancement of details.
5.3 Qualitative Evaluation
Fig. 5 and Fig. 6 show the visual comparison with other state-of-the-art methods on the SID-Sony [4], ELD-Sony [32], LRD [38] and DRV [5] datasets. Images captured in extremely low-light conditions require significant amplification ratios ($\times$300 or -3 EV). These amplified images are plagued by acute noise and color distortion, making most details challenging to discern with the human eye. Compared with previous works, the proposed method shows notable superiority in recovering structural information and enhancing details in noisy regions. In the $\text{2}^{\text{nd}}$ row of Fig. 5, the results produced by previous works tend to be over-smooth due to the excessive denoising while the proposed method present sharp and clean leaves without color bias. In the $\text{5}^{\text{nd}}$ row of Fig. 5, previous works struggle to produce meaningful text information when portions of it are obscured by intense noise in the low-light image. The proposed method, which leverages the potent generative prior from pre-trained Stable Diffusion, exhibits a significantly improved reconstruction of the text.
Thanks to the noise-free generative prior in the decoder $\mathcal{D}$, the sRGB images produced by our proposed method consistently present clean content, irrespective of the degree of noise in low-light inputs.
5.4 Ablation Study
Table 2: The ablation study of the proposed method on ELD-Sony dataset [32]. Different taming strategies (rescaling and DWT) and taming targets (LDM and decoder $\mathcal{D}$) are combined in this study. The best score is denoted as boldfaced.
LDM
Decoder $\mathcal{D}$
PSNR $\uparrow$
SSIM $\uparrow$
LPIPS $\downarrow$
Rescaling - 26.3223 0.7544 0.5423
27.3342 0.8166 0.2934
DWT - 26.9322 0.7652 0.4904
29.8284 0.8906 0.2035
Effectiveness of the DWT
As we mentioned before, 2D discrete wavelet transforms (DWT) not only extract sub-bands for the taming but also downsample the input information to align with the latent size in LDM. Instead of DWT, there are two other intuitive ways for downsampling: rescaling and CNN-based downsampling. In this study, we forego discussion on CNN-based downsampling due to its substantial computational expense. As shown in Fig. 4, simply rescaling the input (e.g., using Bicubic Interpolation) results in the loss of many details and causes the small-scale input to be overwhelmed by noise. Taming the LDM with such input could result in an ambiguous global structure, potentially leading to low-quality sRGB images after decoding. A qualitative comparison between rescaling and DWT taming strategies is shown in Fig. 7.
About taming
The taming modules of the LDM hold significant importance as they steer the generating process toward generating specified latent representations. There is no need to conduct an ablation study on whether taming the LDM since the LDM without taming generates unexpected latent representation which is not related to the low-light input.
The fact that motivates us to tame the decoder $\mathcal{D}$ is that although the low-frequency sub-band (LL) helps the proposed method to tame the LDM for generating a high-correlation latent representation, the sRGB image decoded from this representation by a decoder lacking taming modules, shows much over-generated and irrelevant details.
Fig. 7 and Table 2 show that the decoder $\mathcal{D}$ when utilizing our DWT taming strategy, exhibits substantial improvements in color correction and high-frequency detail maintenance performance.
6 Conclusion
To improve the LLIE performance in extremely dark and noisy images, this work extricates itself from collecting real-world data or carefully designing noise models for synthesizing challenging data. Instead, we propose to tame a pre-trained diffusion model and explore its generative priors for achieving the LLIE task. Our designed taming modules are fed with the low-light input information, and modulate intermediate features within the diffusion model, which allows us to steer the generating process. With the observation of dedicated generative priors in different portions of the diffusion model, we propose to utilize 2D discrete wavelet transforms to divide the LLIE task into two parts: the low-frequency content generation and the high-frequency detail maintenance. Benefiting from the abundant and dedicated generative priors in the diffusion model, extensive experiments demonstrate that taming the diffusion model allows us to achieve the optimized balance between image fidelity and perceptual quality, leading to state-of-the-art LLIE performance.